Find the product of the cube of (-4/3) and square of (-3/2). : Find the product of the cube of (-4/3) and square of (-3/2).
Answer by Alan3354(1216)

Find the product of the cube of (-4/3) and square of (-3/2).
(-4/3)^3 = -64/27
(-3/2)^2 = 9/4
(-64/27)*(9/4) = -576/108
=-16/3
=-5 1/3